Please help me with this: -5(x+4)-2x-6
Answer by LinnW(1048)

distribution rule is a(b + c) = a*b + a*c
for -5(x+4), a = -5 , b = x , c = 4
substituting,
-5(x) + -5(4)
-5x -20
substitute -5x -20 in -5(x+4)-2x-6
-5x -20 -2x-6
combine like terms
-7x -26